OpenStudy (anonymous):

Solve the equation by factoring. 12x^2 + 12x = 24

OpenStudy (anonymous):

can't seem to get this one!!! ugh

OpenStudy (saifoo.khan):

wanna learn the easy way?

OpenStudy (anonymous):

Try factoring out the 12 first.

OpenStudy (anonymous):

YUP!

OpenStudy (saifoo.khan):

12x^2 + 12x - 24 = 0 factor 12 12(x^2 + x -2) = 0 x^ + x - 2 Can u solve it now?

OpenStudy (anonymous):

this is what I have so far 12x^2+12x-24=0 (24x-1) ????

OpenStudy (anonymous):

that's where I get lost!

OpenStudy (anonymous):

When you get to \(x^2 + x -2\), you need to find 2 numbers which will multiply to get \(-2\), and which will add to get \(+1\) (since +x means +1x). Which two numbers would you use?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

You need two numbers. They have to multiply to get \(-2\), which means they are factors of -2. Which numbers are they?

OpenStudy (anonymous):

2 and 1?

OpenStudy (anonymous):

Those are the right numbers, though one of them has to be negative. So it's either \((-2\times +1)\) or \((+2\times -1)\). The two numbers also have to add to get \(+1\). Which of these sets will add together to get \(+1\)?

OpenStudy (anonymous):

(+2x-1)?

OpenStudy (anonymous):

Exactly. That means that your factors are \((x+2)\), and \((x-1)\)

OpenStudy (mimi_x3):

use the quadratic formula its easier